# backgroundcolor=, commentstyle=, breakatwhitespace=false, keywordstyle=, language=Java, stringstyle=, basicstyle=, showstringspaces=false

### v1

@@ 3 Proposed Methodology
  • We reused the benchmark dataset proposed by Maldonado et al. [42] to train an SATD classifier.
  • The pre-trained SATD classifier was then employed to detect SATD comments from open-source projects within the Stack corpus [29].
@@
Tesoro facilitates the detection of technical debt (TD) not only in comments but also through an additional feature: the source code. For SATD tasks, Tesoro offers additional code context, rather than relying solely on comments as in previous studies [42, 16, 58]. To construct the dataset for TD detection in source code, we employed information from SATD comments to identify specific categories of debt within the code. While in prior work the identification of TD was based solely on comments, in this work we identify where TD appears in the source code, without accompanying comments.
@@ 3.1 Source Data
We initially opted for GitHub as our data retrieval source. However, due to rate limit constraints of the GitHub API, [^2] https://docs.github.com/en/rest?apiVersion=2022-11-28 we adopted an alternative dataset: The Stack [29], which has been acknowledged as the most extensive publicly available source code dataset, boasting a permissive license and a substantial size of 3TB. The Stack contains samples from 358 programming languages. In this work, we focus on detecting TD in Java source code. This subset of The Stack yielded a dataset of 26M raw files. Due to constraints in storage and computational resources associated with the code parser and SATD Detection tools, we restricted ourselves to analyzing just 2M of these files. Table I shows the statistics for the dataset across each phase.
TABLE I: Input data information across phases.
@@
3.2 Code Parser Tool
Figure 2: Extraction of comments and functions.
As our emphasis is on detecting TD at the function level, we needed to parse code files from The Stack into individual functions. Since comments serve as the primary annotations for identifying TD within a function, we extracted a collection of functions from each file, each containing a set of comments. We leverage Toolkit [^3] https://github.com/FSoft-AI4Code/CodeText-parser—a tool introduced in our previous work [43], which relies on Tree-sitter [^4] https://tree-sitter.github.io/tree-sitter/—to parse source code into Abstract Syntax Tree (AST) representations, enabling the extraction of functions. Subsequently, we extracted a series of comments associated with each function.
The comments extracted from each function are then annotated following previous research on SATD detection [42, 16, 60, 37]. The labeled information of comments then serves as the ground truth for assigning TD in the function after removing all comments.
3.3 SATD Detection Tool
Since there is a large number of functions and comments, the annotation process requires considerable human labor. Thus, it becomes crucial to choose a subset for annotation purposes. Previous studies [42, 16] showed that the majority of extracted comments do not include TD, with over 90% of comments not implying TD. Therefore, randomly selecting examples for annotation might yield numerous comments that do not contribute to the TD identification process.
Identify comments containing TD: To address this challenge, we developed a TD detection tool to identify comments containing TD within the corpus from Section 3.2. In particular, we constructed a neural model to determine whether a comment contains TD or not. In fact, there have been various SATD detection techniques [18, 55], but we decided to develop a tailored tool on top of pre-trained models as a means to validate thei effectiveness in detecting SATD. The detection tool is a binary classifier, in which all comments containing TD are classified into the positive class, while the rest are assigned to the negative class. Since comments are predominantly in natural language text format, we built the tool using the RoBERTa architecture [39]. Since the tool needed to process a substantial volume of comments, we employed the base version of the model, with 128 million parameters, to balance performance and speed. The Maldonado-62K [16] dataset was leveraged for fine-tuning. The dataset includes more than 62,000 comments, of which 6.5% were identified as containing debt, and categorized into 5 different types of TD, and the majority were identified as non-SATD. We grouped the five classes into the positive class to train the neural classifier using binary cross-entropy loss.
[…]
Detecting TD types of comments: After acquiring the candidate comments, we classified them into five types (design, implementation, defect, test, and documentation) following da Silva Maldonado el at. [16]. Initially, we intend to employ this information to guide annotators, thereby mitigating their workload. However, this information may have biased the annotators. Hence, we utilized this information to investigate TD types that are frequently misunderstood by the model’s capabilities, thus pinpointing examples worthy of annotation (Section 3.4). Instead of employing multiclass classification to detect TD types within comments, we constructed a binary classifier for each type. For instance, when considering TD types $X$, Classifier-$X$ is developed to distinguish whether a comment contains TD type $X$ or not. Similar to identifying comments containing TD, we designate training examples containing TD type $X$ as the positive class, while the remainder are categorized as the negative class. Consequently, we created five classifiers and each of the 1.6 million extracted comments was analyzed by these five classifiers to obtain pseudo-categories. Since these classifiers work independently, a comment can be categorized into more than one class. Figure 3 depicts the overlap categories predicted within a single comment, demonstrating the similarity between the two types of TD. It is shown that design and implementation are the two categories most often confused, a finding consistent with prior studies [37] that tried to merge these categories. However, we maintained these categories separately for more fine-grained evaluation in our dataset.
Figure 3: Overlap categories ratio from multiple binary classifiers prediction on a comment.
3.4 Sampling Strategy
@@
Our objective is not only the selection of examples to explore the detection of TD at the function level but also the identification of comments that could enrich existing datasets. Following existing literature [57, 13, 9, 61], we utilized uncertainty scores to identify challenging instances for annotation, employing the Entropy score. Consider an instance $x$, a model with parameter $\theta$ and $C$ classes, the uncertainty score of the model on sample $x$ is as follows:
$\text{entropy}_{\theta}(x)=-\sum_{i=1}^{C}P_{\theta}(y=i|x)\log(P_{\theta}(y=i|x))$ (1)
@@
     $n$Number of samples for selection.
[…]
     $\mathbf{N}$No. comments extracted from detection tool.
     $\mathbf{D}$Extracted triplet (comment, function, category prediction list) set: $\{d_{i}=(c_{i},f_{i},P_{i})|$ $i=\overline{1,N}\}$.
@@
$\mathbf{Q}\leftarrow\{d_{i}|$ $d_{i}\in\mathbf{D},|P_{i}|>1\}$triangleright Comments that are predicted to contain more than 2 types of TD.
@@
     $\theta\leftarrow\theta_{\mathbf{M}[p]}$
     $s\leftarrow\text{entropy}_{\theta}(c_{j})$triangleright Calculate using Equation 1
     $\mathtt{UnSc}$.add($s$)
@@
$\mathbf{\hat{Q}}\leftarrow\{d_{j}|$ $d_{j}\in\mathbf{D}\backslash\mathbf{Q},j\in\mathtt{top}_{|\mathbf{Q}|}(\mathtt{sID})\}$
@@
3.5 Data Annotation Process
3.5.1 Annotation Group
@@
3.5.2 Labeling process
TABLE II: Annotation Assessment.
Each annotator is provided detailed guidelines that serve as a reference throughout the annotation process. These guidelines include standardized procedures, examples of correctly labeled data, and common pitfalls to avoid. By adhering to these guidelines, the annotators better maintain consistency and reliability across the dataset. Following da Silva Maldonado et al. [42], we developed a tool for the labeling process. However, diverging from the conventional approach of displaying only comments, we also included the corresponding code function as a reference for annotators. Annotators were asked to review both the comments and the corresponding code for labeling. Moreover, we limited the labeling process to our five specific TD types: design, defect, documentation, implementation, and test debts, in addition to non-SATD. The students were given comment and corresponding code, and they had to read both and make a decision. In case, there is no TD contained in the code, so the corresponding label non-SATD was given. Additionally, we conducted Cross-checking and Label Auditing to enhance the quality of the labeling process. Specifically, regarding a data sample, the labeling process is outlined as follows.
  1. 1.
@@ 3.5.2 Labeling process
As shown in Table II, the labeling process was conducted in two phases. In the first phase, 1,400 comments were selected for annotation by seven annotators. Aiming for reliability, every comment was labelled by two students, i.e., each annotator was assigned 400 comments. This phase helps the students familiarize themselves with the labeling task, and establish uniform conventions for the labeling process. Subsequently, in the second phase, there were 3,680 different comments, and each of them was independently evaluated by two students. This resulted in a total of 7,360 comments for the labeling process. To guarantee the reliability of the labeling process, we assessed it using two consensus metrics: Raw Agreement and Inter-Annotator Agreement.
  1. 1.
    Raw Agreement: refers to the count of items for which both annotators assign identical labels, expressed as a percentage of the total items annotated [4].
  2. 2.
    Inter-Annotator Agreement (IAA): Cohen’s Kappa coefficient [32, 19] was applied to quantify the agreement or consistency between different annotators.
Table II presents the annotation scores across two phases. In the initial phase, as the annotators were becoming acquainted with the task, there is a relatively low agreement of 37%, referring to a Fair agreement. However, following reviews and discussions, the agreement strength improved significantly. In the second phase, the Raw Agreement increased by over 35%, and the IAA improved by over 8%, resulting in a Moderate agreement strength. This demonstrates the reliability of our labeling process and the overall quality of the dataset. Following the labeling of TD types for the comments, we aligned the labeled comments with their corresponding code functions, thus obtaining multi-TD type information for an entire code function.
4 Data Characteristics
TABLE III: The comparison between popular SATD benchmarks and Tesoro. NL-sample refers to data in natural language text format, such as comments, pull requests, issues, and commit messages.
Dataset #Code-sample #NL-sample % TD NL-samples #Repo Maldonado-62K [16] - 62,566 6.5 10 4Source-SATD [37] - 95,455 8.5 103 SATD in R [58] - 146,583 3.4 503 Tesoro 1,255 4,981 31.1 974
@@ 4 Data Characteristics
4.1 Dataset for SATD-Related Tasks
We created Tesorocomment, where comments serve as the input source, to support SATD-related tasks, including the identification, classification, and detection of TD. These tasks are structured as multi-class classification problems, with details in Section 5.1. In addition, unlike existing datasets, each comment is associated with its corresponding code, providing a richer context for investigation and analysis.
Figure 4: Category distribution in Tesorocomment. Left: distribution of TD categories within comments containing SATD. Right: percentage of comments that contain versus those that do not contain SATD.
Figure 5: Statistics of Tesorocode. Left: Distribution of the number of comments per function. Right: Distribution of the number of TD types within a function.
Figure 4 presents the statistics of the Tesorocomment dataset, which contains 5,000 labeled comments across six categories: the five TD types and non-SATD. Consistent with previous studies [42, 16, 37], design and implementation debts constitute the majority, with fewer entries for test and documentation, reflecting the real-world distribution. On the other hand, comments with SATD represent a significant portion of the dataset, i.e., 31.1%, which is considerably higher than the proportions in previous datasets (Table III). This underscores the effectiveness of our SATD detection tool in identifying SATD comments, which helps to mitigate the imbalance between SATD and non-SATD comments. Table III shows that our dataset is sourced from 974 repositories. Compared to existing studies Tesorocomment is derived from a more diverse range of sources, capturing a wider variety of commenting and coding styles.
4.2 Dataset for TD detection in Source Code
Next we introduce a dataset named Tesorocode, to support detecting technical debt in source code without relying on natural language comments. Unlike comments, a function can contain multiple types of technical debt; hence, we formulate this scenario as a multi-label classification problem. Specifically, we exclude comments within the function and consider TD types that indicate intrinsic issues in the source code: design, implementation, defect, and test. As a result, Tesorocode presents a challenge for detecting these types of TD within a code function. Table III highlights that no existing dataset has addressed this crucial scenario, underscoring the significance of Tesorocode.
@@ 4.2 Dataset for TD detection in Source Code
5 Experimental Results
@@
We now discuss the experimental results and answer the research questions outlined in Section 1. For RQ1 we first demonstrate that the enhanced dataset improves the performance of existing SATD classification models. RQ2 seeks to assess the significance of source code in detecting SATD comments. We use the function containing the comments as supplementary features to aid in identifying SATD comments. Lastly, we assess the effectiveness of various models in identifying technical debts from source code without relying on SATD comments. In RQ3 our goal is to utilize the curated dataset to detect the presence of different types of technical debts within source code.
TABLE IV: The performance (F1-score) of five PLMs across three tasks when trained on the Maldonado-62K dataset and further enhanced with the additional Tesoro dataset.
Model Identification Classification Detection Maldonado-62K +Tesoro ($\Delta$) Maldonado-62K +Tesoro ($\Delta$) Maldonado-62K +Tesoro ($\Delta$) BERT 87.96 88.99 ( $\uparrow$1.17%) 51.42 55.10 ( $\uparrow$7.16%) 45.99 49.64 ( $\uparrow$7.94%) RoBERTa 89.06 89.96 ( $\uparrow$1.01%) 53.91 55.32 ( $\uparrow$2.62%) 46.13 52.86 ( $\uparrow$14.59%) UniXCoder 88.38 88.42 ($\uparrow$0.05%) 54.82 54.99 ( $\uparrow$0.31%) 50.94 52.11 ( $\uparrow$2.30%) CodeBERT 88.74 90.06 ( $\uparrow$1.49%) 57.50 63.70 ( $\uparrow$10.78%) 53.79 55.60 ( $\uparrow$3.36%) GraphCodeBERT 89.94 90.12 ( $\uparrow$0.20%) 58.00 60.44 ( $\uparrow$4.21%) 49.12 56.87 ( $\uparrow$15.78%)
5.1 RQ1: Do the manually classified comments contribute to an improvement in the detection of SATD?
Motivation: Our primary focus in this RQ is to assess whether the newly identified SATD comments from the Stack corpus contribute positively to the performance of SATD comment detection. Our objective is to enhance the state-of-the-art benchmark dataset introduced by da Silva Maldonado et al. [16], henceforth referred to as Maldonado-62K. This dataset comprises 62,566 comments, of which 4,071 (approximately 6.5%) encompass one of five categories of technical debt. While the dataset’s volume is relatively limited, making it challenging to apply deep learning models, it is crucial to increase the diversity of data samples within each category to enhance the training performance of SATD detection models.
Approach: For this RQ, our objective is to examine whether manually classified comments from the Tesorocomment dataset improve the detection of SATD comments. Following previous studies [16, 53, 55, 59], we employed a cross-project experimental approach. The benchmark dataset Maldonado-62K comprises 10 projects, divided into 10 validation folds. We used one fold (one project) for testing and the remaining nine folds (nine other projects) for training. To prevent data leakage, we removed duplicate entries from the entire dataset before splitting it, resulting in 38,269 samples. We then analyzed the impact of incorporating Tesoro during training, denoted as $+$Tesoro, compared to using only the state-of-the-art dataset, Maldonado-62K, across three scenarios: SATD-Identification, SATD-Classification and SATD-Detection.
  1. 1.
  2. 2.
  3. 3.
For each scenario, we used the same test set while training the model separately with the Maldonado-62K and with $+$Tesoro datasets. Research in the field of SATD detection has highlighted the promising results of PLMs [18, 58, 59], thus we also employed these models to detect SATD comments. Specifically, we experimented with BERT [17], RoBERTa [39], CodeBERT [22], UniXCoder [23] and GraphCodeBERT [24]. PLMs are employed as comment encoders, followed by a fully connected network dedicated to downstream tasks such as identification, classification, and detection, as previously described. All models are fine-tuned for 10 epochs using a batch size of 32 and a learning rate of $51e-$ 5. Given the significant class imbalance in the dataset, the F1 score to was used to evaluate performance.
Result: Table IV depicts the performance of five different PLMs across three aforementioned scenarios $\mathbb{S}_{1}$, $\mathbb{S}_{2}$ and $\mathbb{S}_{3}$. We see that incorporating Tesoro during the training phase consistently boosts the performance of all PLMs across all tasks. Specifically, the identification of SATD comments from the entire set of comments demonstrates an improvement ranging from 0.05% to 1.49% across all models when the training phase is supplemented with the Tesoro dataset, as opposed to relying solely on the Maldonado-62K dataset. In this scenario, the dataset demonstrates a considerable imbalance, with non-SATD data making up more than 90% of the total. The improvement highlights the effectiveness of the Tesoro comment in mitigating the data imbalance issue, and this is significant because real-world scenarios will have a similar imbalance. In the context of SATD detection, BERT and RoBERTa exhibited lower performance compared to the other three code-based PLMs. However, training with $+$Tesoro considerably improves the performance of these two models, leading to comparable results across all the models. For instance, applying $+$Tesoro during training improved BERT’s performance by 7.94% and RoBERTa’s by 14.59% in detecting various types of SATD comments. A similar trend is observed in other scenarios, where all models show enhanced performance when utilizing $+$Tesoro. Specifically, the performance in classifying the five SATD types increases by approximately 0.31% to 10.78% for all PLM models. Additionally, the results indicate that CodeBERT and its variant, GraphCodeBERT, achieve the highest performance across all scenarios, highlighting the advantage of code-based PLMs for the detection of SATD comments.
Figure 6: An in-depth analysis of CodeBERT and RoBERTa performance across three scenarios for 10 projects.
Finally, we conducted an in-depth analysis of the results from 10 large-scale open source projects across all three scenarios. In this setup, each project serves as a test fold, while the PLMs are trained on the remaining 9 projects. For illustrative purposes, we present the findings for CodeBERT and RoBERTa, as depicted in Figure 6. Overall, it is evident that training the models with $+$Tesoro results in improved performance across all projects. In particular, in the $\mathbb{S}_{1}$ task, the performance improved across all tested projects for both models, highlighting the enhanced training set’s effectiveness in addressing the imbalance issue in the original dataset. Furthermore, 7 out of the 10 folds show an increase in F1-score when training CodeBERT on classification and detection tasks with $+$Tesoro, while RoBERTa exhibits significant improvement in 9 out of 10 test sets for the SATD detection task.
Answer to RQ1 There is an improvement in the prediction performance when Tesoro is incorporated into the training, validating the efficacy of our data pipeline in selecting informative samples, and proving a robust annotation process. CodeBERT and GraphCodeBERT consistently achieve superior performance and notable improvements across all three tasks, highlighting the advantages of employing code-based PLMs for SATD comment detection.
@@ 5.1 RQ1: Do the manually classified comments contribute to an improvement in the detection of SATD?
5.2 RQ2: Does the inclusion of source code help to enhance the detection of technical debt?

### v2

@@ 3. Proposed Methodology
  • We reused the benchmark dataset proposed by Maldonado et al. (Maldonado and Shihab, 2015) to train an SATD classifier.
  • The pre-trained SATD classifier was then employed to detect SATD comments from open-source projects within the Stack corpus (Kocetkov et al., 2022).
@@
Tesoro facilitates the detection of technical debt (TD) not only in comments but also through an additional feature: the source code. For SATD tasks, Tesoro offers additional code context, rather than relying solely on comments as in previous studies (Maldonado and Shihab, 2015; da Silva Maldonado et al., 2017; Sharma et al., 2022). To construct the dataset for TD detection in source code, we employed information from SATD comments to identify specific categories of debt within the code. While in prior work the identification of TD was based solely on comments, in this work we identify where TD appears in the source code, without accompanying comments.
@@ 3.1. Source Data
We initially opted for GitHub as our data retrieval source. However, due to rate limit constraints of the GitHub API [^2] https://docs.github.com/en/rest?apiVersion=2022-11-28, we adopted an alternative dataset: The Stack (Kocetkov et al., 2022), which has been acknowledged as the most extensive publicly available source code dataset, boasting a permissive license and a substantial size of 3TB. The Stack contains samples from 358 programming languages. In this work, we focus on detecting TD in Java source code. This subset of The Stack yielded a dataset of 26M raw files. Due to constraints in storage and computational resources associated with the code parser and SATD Detection tools, we restricted ourselves to analyzing just 2M of these files. Table 1 shows the statistics for the dataset across each phase.
3.2. Code Parser Tool
As our emphasis is on detecting TD at the function level, we needed to parse code files from The Stack into individual functions. Since comments serve as the primary annotations for identifying TD within a function, we extracted a collection of functions from each file, each containing a set of comments. We leverage Toolkit [^3] https://github.com/FSoft-AI4Code/CodeText-parser—a tool introduced in our previous work (Manh et al., 2023), which relies on Tree-sitter [^4] https://tree-sitter.github.io/tree-sitter/—to parse source code into Abstract Syntax Tree (AST) representations, enabling the extraction of functions. Subsequently, we extracted a series of comments associated with each function.
The comments extracted from each function are then annotated following previous research on SATD detection (Maldonado and Shihab, 2015; da Silva Maldonado et al., 2017; Sierra et al., 2019; Li et al., 2023). The labeled information of comments then serves as the ground truth for assigning TD in the function after removing all comments.
Figure 2. Extraction of comments and functions.
3.3. SATD Detection Tool
Since there is a large number of functions and comments, the annotation process requires considerable human labor. Thus, it becomes crucial to choose a subset for annotation purposes. Previous studies (Maldonado and Shihab, 2015; da Silva Maldonado et al., 2017) showed that the majority of extracted comments do not include TD, with over 90% of comments not implying TD. Therefore, randomly selecting examples for annotation might yield numerous comments that do not contribute to the TD identification process.
Identify comments containing TD: To address this challenge, we developed a TD detection tool to identify comments containing TD within the corpus from Section 3.2. In particular, we constructed a neural model to determine whether a comment contains TD or not. In fact, there have been various SATD detection techniques (Di Salle et al., 2022; Sala et al., 2021), but we decided to develop a tailored tool on top of pre-trained models as a means to validate thei effectiveness in detecting SATD. The detection tool is a binary classifier, in which all comments containing TD are classified into the positive class, while the rest are assigned to the negative class. Since comments are predominantly in natural language text format, we built the tool using the RoBERTa architecture (Liu et al., 2019). Since the tool needed to process a substantial volume of comments, we employed the base version of the model, with 128 million parameters, to balance performance and speed. The Maldonado-62K (da Silva Maldonado et al., 2017) dataset was leveraged for fine-tuning. The dataset includes more than 62,000 comments, of which 6.5% were identified as containing debt, and categorized into 5 different types of TD, and the majority were identified as non-SATD. We grouped the five classes into the positive class to train the neural classifier using binary cross-entropy loss.
[…]
Detecting TD types of comments: After acquiring the candidate comments, we classified them into five types (design, implementation, defect, test, and documentation) following da Silva Maldonado el at. (da Silva Maldonado et al., 2017). Initially, we intend to employ this information to guide annotators, thereby mitigating their workload. However, this information may have biased the annotators. Hence, we utilized this information to investigate TD types that are frequently misunderstood by the model’s capabilities, thus pinpointing examples worthy of annotation (Section 3.4). Instead of employing multiclass classification to detect TD types within comments, we constructed a binary classifier for each type. For instance, when considering TD types $X$, Classifier-$X$ is developed to distinguish whether a comment contains TD type $X$ or not. Similar to identifying comments containing TD, we designate training examples containing TD type $X$ as the positive class, while the remainder are categorized as the negative class. Consequently, we created five classifiers and each of the 1.6 million extracted comments was analyzed by these five classifiers to obtain pseudo-categories. Since these classifiers work independently, a comment can be categorized into more than one class. Figure 3 depicts the overlap categories predicted within a single comment, demonstrating the similarity between the two types of TD. It is shown that design and implementation are the two categories most often confused, a finding consistent with prior studies (Li et al., 2023) that tried to merge these categories. However, we maintained these categories separately for more fine-grained evaluation in our dataset.
Table 1. Input data information across phases.
@@ 3.3. SATD Detection Tool
Table 2. Annotation Assessment.
Figure 3. Overlap categories ratio from multiple binary classifiers prediction on a comment.
3.4. Sampling Strategy
@@
Our objective is not only the selection of examples to explore the detection of TD at the function level but also the identification of comments that could enrich existing datasets. Following existing literature (Settles, 2009; Coleman et al., 2019; Chitta et al., 2021; Sorscher et al., 2022), we utilized uncertainty scores to identify challenging instances for annotation, employing the Entropy score. Consider an instance $x$, a model with parameter $\theta$ and $C$ classes, the uncertainty score of the model on sample $x$ is as follows:
(1) $\text{entropy}_{\theta}(x)=-\sum_{i=1}^{C}P_{\theta}(y=i|x)\log(P_{\theta}(y=i|x))$
@@
  nnNumber of samples for selection.
  ccList of TD types.
  NNNo. comments extracted from detection tool.
  DDExtracted triplet (comment, function, category prediction list) set: $\{d_{i}=(c_{i},f_{i},P_{i})|$ $i=\overline{1,N}\}$.
@@
$Q\leftarrow\{d_{i}|$ $d_{i}\in D,|P_{i}|>1\}$ $\triangleright$ Comments that are predicted to contain more than 2 types of TD.
@@
  $\theta\leftarrow\theta_{M[p]}$
  $s\leftarrow\text{entropy}_{\theta}(c_{j})$ $\triangleright$ Calculate using Equation 1
  UnScUnSc.add($s$)
@@
return $D$
3.5. Data Annotation Process
3.5.1. Annotation Group
@@
3.5.2. Labeling process
Each annotator is provided detailed guidelines that serve as a reference throughout the annotation process. These guidelines include standardized procedures, examples of correctly labeled data, and common pitfalls to avoid. By adhering to these guidelines, the annotators better maintain consistency and reliability across the dataset. Following da Silva Maldonado et al. (Maldonado and Shihab, 2015), we developed a tool for the labeling process. However, diverging from the conventional approach of displaying only comments, we also included the corresponding code function as a reference for annotators. Annotators were asked to review both the comments and the corresponding code for labeling. Moreover, we limited the labeling process to our five specific TD types: design, defect, documentation, implementation, and test debts, in addition to non-SATD. The students were given comment and corresponding code, and they had to read both and make a decision. In case, there is no TD contained in the code, so the corresponding label non-SATD was given. Additionally, we conducted Cross-checking and Label Auditing to enhance the quality of the labeling process. Specifically, regarding a data sample, the labeling process is outlined as follows.
  1. (1)
@@ 3.5.2. Labeling process
As shown in Table 2, the labeling process was conducted in two phases. In the first phase, 1,400 comments were selected for annotation by seven annotators. Aiming for reliability, every comment was labelled by two students, i.e., each annotator was assigned 400 comments. This phase helps the students familiarize themselves with the labeling task, and establish uniform conventions for the labeling process. Subsequently, in the second phase, there were 3,680 different comments, and each of them was independently evaluated by two students. This resulted in a total of 7,360 comments for the labeling process. To guarantee the reliability of the labeling process, we assessed it using two consensus metrics: Raw Agreement and Inter-Annotator Agreement.
  1. (1)
    Raw Agreement: refers to the count of items for which both annotators assign identical labels, expressed as a percentage of the total items annotated (Artstein, 2017).
  2. (2)
    Inter-Annotator Agreement (IAA): Cohen’s Kappa coefficient (Landis and Koch, 1977; Emam, 1999) was applied to quantify the agreement or consistency between different annotators.
Table 2 presents the annotation scores across two phases. In the initial phase, as the annotators were becoming acquainted with the task, there is a relatively low agreement of 37%, referring to a Fair agreement. However, following reviews and discussions, the agreement strength improved significantly. In the second phase, the Raw Agreement increased by over 35%, and the IAA improved by over 8%, resulting in a Moderate agreement strength. This demonstrates the reliability of our labeling process and the overall quality of the dataset. Following the labeling of TD types for the comments, we aligned the labeled comments with their corresponding code functions, thus obtaining multi-TD type information for an entire code function.
4. Data Characteristics
Table 3. The comparison between popular SATD benchmarks and Tesoro. NL-sample refers to data in natural language text format, such as comments, pull requests, issues, and commit messages.
Dataset #Code-sample #NL-sample % TD NL-samples #Repo
Maldonado-62K (da Silva Maldonado et al., 2017) - 62,566 6.5 10
4Source-SATD (Li et al., 2023) - 95,455 8.5 103
SATD in R (Sharma et al., 2022) - 146,583 3.4 503
Tesoro 1,255 4,981 31.1 974
@@ 4. Data Characteristics
4.1. Dataset for TD detection in Source Code
We introduce a dataset named Tesorocode, to support detecting technical debt in source code without relying on natural language comments. Unlike comments, a function can contain multiple types of technical debt; hence, we formulate this scenario as a multi-label classification problem. Specifically, we exclude comments within the function and consider TD types that indicate intrinsic issues in the source code: design, implementation, defect, and test. As a result, Tesorocode presents a challenge for detecting these types of TD within a code function. Table 3 highlights that no existing dataset has addressed this crucial scenario, underscoring the significance of Tesorocode.
@@ 4.1. Dataset for TD detection in Source Code
4.2. Dataset for SATD-Related Tasks
Since we construct Tesorocode using information from comments and later exclude these comments, preserving them can help support existing SATD datasets. Therefore, we created Tesorocomment, where comments serve as the input source, to support SATD-related tasks, including the identification, classification, and detection of TD. These tasks are structured as multi-class classification problems, with details in Section 5.1. In addition, unlike existing datasets (Table 3), each comment in Tesorocomment is associated with its corresponding code, providing a richer context for investigation and analysis.
Figure 4. Category distribution in Tesorocomment. Left: distribution of TD categories within comments containing SATD. Right: percentage of comments that contain versus those that do not contain SATD.
Figure 4 presents the statistics of the Tesorocomment dataset, which contains 5,000 labeled comments across six categories: the five TD types and non-SATD. Consistent with previous studies (Maldonado and Shihab, 2015; da Silva Maldonado et al., 2017; Li et al., 2023), design and implementation debts constitute the majority, with fewer entries for test and documentation, reflecting the real-world distribution. On the other hand, comments with SATD represent a significant portion of the dataset, i.e., 31.1%, which is considerably higher than the proportions in previous datasets (Table 3). This underscores the effectiveness of our SATD detection tool in identifying SATD comments, which helps to mitigate the imbalance between SATD and non-SATD comments. Table 3 shows that our dataset is sourced from 974 repositories. Compared to existing studies Tesorocomment is derived from a more diverse range of sources, capturing a wider variety of commenting and coding styles.
Figure 5. Statistics of Tesorocode. Left: Distribution of the number of comments per function. Right: Distribution of the number of TD types within a function.
5. Experimental Results
@@
We now discuss the experimental results and answer the research questions outlined in Section 1. RQ1 seeks to assess the significance of source code in detecting SATD comments. We use the function containing the comments as supplementary features to aid in identifying SATD comments. Lastly, we assess the effectiveness of various models in identifying technical debts from source code without relying on SATD comments. In RQ2 our goal is to utilize the curated dataset to detect the presence of different types of technical debts within source code. For RQ3 we demonstrate that the enhanced dataset improves the performance of existing SATD classification models.
@@
5.1. RQ1: Does the inclusion of source code help to enhance the detection of technical debt?